OpenStudy (anonymous):

Factor completely. x^3 + x^2 - 25x - 25

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow x^2(x + 1) - 25(x + 1) }\) \(\Large \color{MidnightBlue}{\rightarrow (x^2 - 25)(x + 1) }\)

OpenStudy (anonymous):

a. (x+ 1)(x- 5)(x+ 5) b. (x+ 5)(x- 1)(x+ 5) c. (x- 1)(x+ 5)(x- 5)

OpenStudy (anonymous):

(x - 5)(x +1)(x + 5) = 0 x - 5 = 0 => x = 5 x + 1 = 0 => x = -1 x + 5 = 0 => x = -5 Hence, x = 5 or x = -1 or x = -5

OpenStudy (anonymous):

use grouping method : (x^3+x^2)-(25x-25) (x^2-25)(x+1) (x-5)(x+5)(x+1)